FIG. P4.12 4.37. In Problem 36, let m = 0.5 kg, l = 0.9 m, k 3x 103 N/m, c 10 N?and F 5 sin 10t N. Determine the complete solution as a function of time for t following initial conditions 0,-0 and 60 = 3 rad/s. 4.38. Assuming small angular oscillations, derive the differential equation of motion of the vibratory system shown in Fig. P13.

Problem 4.38
Answer #1

Cos From the ameby shown above, asSp iom Fres body daagram Momont of inerわa no

Hence, the equation of motion is ---

\mathbf{m L \ddot{\theta} + c L \dot{\theta} + m g \theta = - c \omega_r Y_0 Cos \omega_r t}